OpenStudy (dottiecat):

Three girls are preparing information packets about an upcoming fundraiser event. Nicole prepared 270 packets in 45 min. Hillary prepared 450 packets in 1.5 hours McKenzie prepared 280 packets in 1 hour and 10 min. a. Who works the fastest? Explain how you solved this problem. (Hint: make sure the units you are comparing are the same) b. Each girl worked another 2 hours at her same rate. Who made the most packets that day? Explain how you solved this problem.

OpenStudy (dottiecat):

@Qwertty123 @Study312

OpenStudy (qwertty123):

Divided the amount of packets by the time. Also turned the time into minutes, not hours.

OpenStudy (dottiecat):

first one is 6. second one is 300

OpenStudy (qwertty123):

What... Um

OpenStudy (dottiecat):

oops hold on

OpenStudy (study312):

@Qwertty123 take the wheel

OpenStudy (dottiecat):

450 / 90 = 5 because 1.5 hours is 90 minutes

OpenStudy (dottiecat):

and 280 / 70 = 4 because 70 = 1 hour + 10 mins

OpenStudy (dottiecat):

right?

OpenStudy (dottiecat):

so for (A) McKenzie is the fastest, right?

OpenStudy (qwertty123):

YES! :D

OpenStudy (dottiecat):

awesome k lets start on letter b

OpenStudy (dottiecat):

oh my. where do we begin o-O

OpenStudy (qwertty123):

Divided! Okay... hold on

OpenStudy (dottiecat):

okay so this is what I think. McKenzie: 1 hr + 10 = 70 mins. 70 mins + 120 mins = 840.. Hillary: 90 mins + 120 mins = 210 minutes. Nicole: 45 mins + 120 mins = 165 mins

OpenStudy (dottiecat):

McKenzie made the most packets?

OpenStudy (dottiecat):

idk if I did that right .-.

OpenStudy (study312):

the first one is already in mins

OpenStudy (dottiecat):

I know that :P

OpenStudy (dottiecat):

Did I do it incorrectly?

OpenStudy (study312):

you want to know how many packets there were each minute. so 270 ÷ 45

OpenStudy (qwertty123):

This should be it McKenzie: 4 packets a min.

OpenStudy (qwertty123):

Did I do something wrong?

OpenStudy (dottiecat):

That's what I got

OpenStudy (study312):

how did you get 4?

OpenStudy (study312):

270 ÷ 45 = 6 450 ÷ 110 = 4 280 ÷ 70 = 4 correct me if i'm wrong

OpenStudy (study312):

your probably correct

OpenStudy (study312):

60 + 50 oooooohhhhhhh

OpenStudy (study312):

i forgot 1.5 hours is 1 and a half hours

OpenStudy (study312):

sorry

OpenStudy (dottiecat):

haha no problem at all. so would I be correct? :o

OpenStudy (study312):

yes